Innovation Company is thinking about marketing a new software product. Upfront costs to market and develop the product are $4.97 million. The product is expected to generate profits of $1.19 million per year for ten years. The company will have to provide product support expected to cost $90,000 per year in perpetuity. Assume all profits and expenses occur at the end of the year a. What is the NPV of this investment if the cost of capital is 5.9%? Should the firm undertake the project? Repeat the analysis for discount rates of 1.0% and 17.8%, respectively b. What is the IRR of this investment opportunity? What does the IRR rule indicate about this investment?
Answer #1

NPV at 5.9% = -Initial cost + PV of annuity – PV of perpetuity

= -4,970,000 + 1190000*(1-1/(1+0.059)^10)/0.059 – 90000/5.9%

=-4970,000+8800149.1 -1525423.7

=2304725.4

NPV at 1%= = -Initial cost + PV of annuity – PV of perpetuity

= -4,970,000 + 1190000*(1-1/(1+0.01)^10)/0.01 – 90000/1%

=-2699148

NPV at 17.8%= = -Initial cost + PV of annuity – PV of perpetuity

= -4,970,000 + 1190000*(1-1/(1+0.178)^10)/0.178 – 90000/17.8%

=-89418.86

IRR is that rate at which NPV = 0

Hence

Initial cost = PV of annuity – PV of perpetuity

4970,000= 1190000*(1-1/(1+IRR)^10)/IRR -90000/IRR

Using trial and error metjhod, we get the IRR as 1.4988% or 1.5%

If the cost of capital is 5.9%, the IRR is much lower and so the project should not be undertaken.